OpenStudy (anonymous):

HELP PLEASE! Which statement best explains the movement of electric current from the clouds to the ground during a lightning storm? A. the ground and the clouds are both negatively charged. B. the ground and the clouds are both positively charged C. the ground is positively charged and the clouds are negatively charged D. the ground is negatively charged and the clouds are neutral ???

OpenStudy (jamesj):

Discharge between any two objects (clouds, ground, my hand, the cat, ...) happens when the potential difference between them is sufficiently large. Which of these four options suggests that the potential difference between the clouds and the ground is the largest?

OpenStudy (anonymous):

idk you tell me ;)

OpenStudy (anonymous):

and the answer is..............

OpenStudy (jamesj):

Make it concrete if it helps you. Let G be the charge of the ground and C the charge of the clouds. In the first scenario, A., both ground and clouds have negative charge, suppose G = -10 Volts and C = -10 V Then the potential difference between them G - C = -10 - (-10) = 0 So that is not conductive to a discharge and a current flowing between them. You analyze now the other three options and figure out which one has the greatest potential difference.

OpenStudy (anonymous):

C.GROUND CARRIES POSITIVE CHARGE AND CLOUDS CARRY POSITIVE

OpenStudy (jamesj):

To begin with: that is the wrong answer. Next, it completely disrespects the process we've got going on here where we are trying to lead the student to the answer, instead of giving it to them without their being able to replicate the answer themselves.

OpenStudy (radar):

Electrons (being themselves negative) travel to a point of positive potential, or to a lesser negative potential. With this in mind you can determine the charge relationship between ground and clouds.